Pulled water to let top off extra. Trimmed echinata, added pieces back to colony. Cyano algae is gone! One remaining bottle of bacteria substrate sauce.
View May 7, 2021 20:53
Skimmer currently off for cyano abatement. To be switched back on next week.
View April 30, 2021 19:46
Feeding regimen changed. 2x2" Reef Frenzy one hour after lights come on (2pm). 2x3" sheet of Nori seaweed (5pm). Four cubes (Mysis Shrimp or Ocean Plankton) a few hours before lights out (8pm).
View April 26, 2021 22:18
Replaced Trident Reagent A. Left sump kept low to top off R/O. Replaced all R/O unit filters. More LRS and Nori (and updated feeding regimen) coming next week.
View April 23, 2021 19:26
Water change deferred to next week because of weather. Scooped out thick cyano algae, stirred dusting.
View April 17, 2021 04:41
Bottle of bacteria/coral food left in basement bar mini fridge for every other week dosing to help with cyano. DO NOT DRINK.
View April 17, 2021 04:41
Discovered that EcoTech battery backup was connected to return pumps (which require more power than the batteries provide). Now connected to right side Vortech powerheads. Tested and working properly.
View April 17, 2021 00:11
Yellow Tang died. Beaten up and had Lateral Line Erosion disease. One of three remains. (Endemic varieties no longer sold from Hawaii.)
View April 15, 2021 02:11
Cyanobacteria infestation increasing. Scooped it out. Skimmer off as a test for a week to see if nitrates increase. Cyano-eating bacteria to be introduced next week. Waiting for Dragon Goby to come back in stock.
View April 9, 2021 19:02
Power lost momentarily three times between 7PM and midnight. Backup batteries for Vortech powerheads did not engage.
View April 7, 2021 01:05